OpenStudy (anonymous):

A puck of mass 1 kg collides with a wall. The puck will explode if hit with a force greater than 15 N. Before the collision the puck's velocity is 10 m/s. After the collision the puck's velocity is -10 m/s. Assume a constant force is applied to the puck during the collision. What is the shortest time interval over which this puck could have been hit such that it did not explode?

OpenStudy (mani_jha):

\[Force=(Change \in momentum)/Time\] If time has to be the shortest, force has to be the largest. Can you calculate the change in momentum?

OpenStudy (anonymous):

how would you find the change in momentum?

OpenStudy (anonymous):

Change in momentum is m(v2-v1) = 1(-10-10) = -20 kg m/s Force max = 15 N = rate of change of momentum = 20/t t = 20/15 = 1.3333 seconds (min)

OpenStudy (anonymous):

thank you!